OpenStudy (anonymous):

Simplify 3/x+1 + 4/x-2 Can someone show me how to solve this?

OpenStudy (anonymous):

solve as in "add"?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[\frac{3}{x+1}+\frac{4}{x-2}=\frac{3(x-2)+4(x+1)}{(x+1)(x-2)}\] then some algebra to clean up the numerator

OpenStudy (anonymous):

What was the first step?

OpenStudy (anonymous):

the first step is to add like you would add two fractions \[\frac{a}{b}+\frac{c}{d}=\frac{ad+bc}{ad}\] in this case \(a=3, b=x+1, c=4, d=x-2\)

OpenStudy (anonymous):

Do I distribute in the numerators?